solving equations and inequalities
x+1/3X-1 + 2X+1/3X-2 = -1

Answer by stanbon(75887)

solving equations and inequalities
x+1/3X-1 + 2X+1/3X-2 = -1
----
Multiply thru by (3X-1)(3X-2) to get:
(x+1)(3x-2) + (2x+1)(3x-1) = -(3x-1)(3x-2)
----------------------------------------------
3x^2 +x -2 + 6x^2 +x -1 = -9x^2 + 9x -2
-------------------------------------------
9x^2 + 2x -3 = -9x^2 + 9x -2
-------
18x^2 - 7x -1 = 0
18x^2 -9x+2x -1 = 0
---
9x(2x-1) + (2x-1) = 0
------
(2x-1)(9x+1) = 0
-------
x = 1/2 or x = -1/9
==========================
Cheers,
Stan H.